KyledaGreat:

g(x) = x^2 + x − 6 Step 2 of 2 : Sketch the graph by identifying where the graph is positive and where the graph is negative. Select the area above the axis to plot the positive curve for that region, and select the area below the axis to plot the negative curve for that region. Use the Reset button to remove all curves.

jhonyy9:

so like the first step try get the zeros of this quadratic

KyledaGreat:

KyledaGreat:

is this right ?

jhonyy9:

what are the roots of this quadratic ?

KyledaGreat:

x = 2, -3

jhonyy9:

do you can factoriz it completly ?

KyledaGreat:

no

jhonyy9:

@kyledagreat wrote:
x = 2, -3
i think not really

jhonyy9:

-3 +2 = -1 not +1

jhonyy9:

@jhonyy9 wrote:
-3 +2 = -1 not +1
for the b value

jhonyy9:

maybe this (x+3)(x-2) = ?

jhonyy9:

this is true ?

KyledaGreat:

yes sir

jhonyy9:

yes so from this result that your graph is right but you need be sure in your calcule

jhonyy9:

step 2 . ???

KyledaGreat:

umm

KyledaGreat:

i don't think so

jhonyy9:

you need select these areas

KyledaGreat:

oh that's true ok

jhonyy9:

Step 2 of 2 : Sketch the graph by identifying where the graph is positive and where the graph is negative.

KyledaGreat:

okay , was it right to have done the graph based on the picture ?

jhonyy9:

yes right but now you need solve the point 2

jhonyy9:

Step 2 of 2 : Sketch the graph by identifying where the graph is positive and where the graph is negative. this ...

jhonyy9:

use your graph so this is easy

KyledaGreat:

how could i do that

jhonyy9:

do you know where the graph is positive ? and where is negativ ?

KyledaGreat:

I don’t believe so

KyledaGreat:

1 attachment
KyledaGreat:

is it like that ?

KyledaGreat:

are any of them right that i just did ?

KyledaGreat:

jhonny ?

jhonyy9:

@kyledagreat wrote:
positive right but negativ not marked right

KyledaGreat:

which one

jhonyy9:

on the first one posted image with vectors

jhonyy9:

the second negativ quadrant is on the right corner down

jhonyy9:

Step 2 of 2 : Sketch the graph by identifying where the graph is positive and where the graph is negative.

KyledaGreat:

jhonyy9:

what you marked there ? positive - negativ parts ???

KyledaGreat:

yes

KyledaGreat:

does it look like it

jhonyy9:

please mark it with plus or minus signs

KyledaGreat:

i'm sorry for the bluriness , it couldn't focus